# Neural Networks Learn Statistics of Increasing Complexity

### v1

@@ 2 Theory and Methods
where $\alpha$ is a multi-index, or a $d$-tuple assigning an integer to each coordinate of $\mathbf{x}$. Recall that taking a vector to the power of a multi-index denotes a product of the components of the vector, where each component of the index indicates the multiplicity: e.g. if $\alpha=(1,4,6)$, the expression $(\mathbf{x}-\boldsymbol{\mu})^{\alpha}$ denotes the product $(x_{1}-\mu_{1})(x_{2}-\mu_{2})^{4}(x_{3}-\mu_{3})^{6}$. Similarly, ($\partial^{\alpha}\mathcal{L}$) is shorthand for the mixed partial derivative $\frac{\partial^{11}}{\partial x_{1}\partial x_{2}^{4}\partial x_{3}^{6}}\mathcal{L}$. The factorial $\alpha!$ denotes the product of the factorials of the components: $1!\times 4!\times 6!=17280$.
@@ 2.3 Optimal Transport
Figure 4: Rows 1-3 show how Gaussian optimal transport affects the example CIFAR-10 airplane, bird and truck images. Each row starts with the original unedited image on the left, with each subsequent column showing the effects of editing that image’s first two moments to match the class-conditional distributions of a particular target class (top). The bottom left image is the average of all test set images. Black borders indicate either an original image, or a “self-to-self” edited image.
@@
Given $k$ image classes, each containing tensors of shape $C\times H\times W$, we unroll the tensors into vectors of size $CHW$, then compute their sample means and covariance matrices, [^7] Because this is a high-dimensional covariance matrix with dimension only 1-3 times smaller than the sample size, we apply the asymptotically optimal linear shrinkage method proposed by Bodnar et al. (2014) to improve our estimate of the population covariance and increase numerical stability. and plug these statistics into Eq. 3 to get the $k(k-1)$ optimal transport maps from each class to every other class.
@@ 2.4 Maximum Entropy Sampling
One problem with using Gaussians to generate synthetic images is that natural images are constrained to a hypercube: RGB pixel intensities are in the range $[0,255]$, but nonsingular Gaussian distributions assign positive probability density to all of $\mathbb{R}^{d}$, so that a typical sample will often lie outside the hypercube of natural images. [^9] Strictly speaking, Gaussians also violate the assumption of compact support that we made earlier. In high dimension, though, almost all the probability mass of a Gaussian is contained in the typical set, a compact region near the boundary of an ellipsoid surrounding the mean (Carpenter, 2017). We might expect neural networks to quickly adapt to such a simple box constraint on the support, so we would ideally like to subject our synthetic images to this constraint.
@@ 2.5 Discrete Domains
[$n$-gram statistics are moments] Let $\mathcal{V}^{N}$ be the set of token sequences of length $N$ drawn from a finite vocabulary $\mathcal{V}$, let $P$ be a distribution on $\mathcal{V}^{N}$, and let $f:\mathcal{V}^{N}\rightarrow\{0,1\}^{N\cdot|\mathcal{V}|}$ be the function that encodes a length-$N$ sequence of tokens as a flattened concatenation of $N$ one-hot vectors of dimension $|\mathcal{V}|$. Let $f_{\sharp}P$ be the pushforward of $P$ through this one-hot encoding, i.e. its analogue in $\{0,1\}^{N\cdot|\mathcal{V}|}$.
@@ 3 Image Classification
Because Gaussian optimal transport (Sec. 2.3) requires $O(d^{3})$ compute and $O(d^{2})$ memory, [^12] The covariance matrix has $d^{2}$ elements, where $d$ is the number of pixels, so it is actually $O(n^{4})$ in the height or width dimension of the image. We also ran into a software limitation in early experiments where NumPy and PyTorch eigensolvers would crash when fed the very large covariance matrices produced by high-resolution image datasets; see PyTorch issue #92141 for discussion. we focus on datasets with $32\times 32$ or $64\times 64$ resolution images for our primary experiments. Specifically, we examine the popular image classification datasets CIFAR-10 (Krizhevsky et al., 2009), Fashion MNIST (Xiao et al., 2017), MNIST (LeCun et al., 1998), and SVHN (Netzer et al., 2011).
We also build a new image classification dataset, CIFARNet, consisting of 200K images at $64\times 64$ resolution sampled from ImageNet-21K, using ten coarse-grained classes that roughly match those of CIFAR-10. The larger number of images per class (20K) allows us to get a good estimate of the class-conditional covariance matrices needed for Gaussian optimal transport, which at this resolution contain $(3\times 64\times 64)^{2}\approx 1.5\times 10^{8}$ entries each. See Appendix B for more details on CIFARNet.
@@ Appendix B CIFARNet dataset
We ensured class balance by randomly sampling 20K images from each synset. Images were directly resized to $64\times 64$ resolution without center cropping.
@@ Appendix D Discrete domain proofs
where $f(x)_{1}^{0}$ denotes the first component of $f(x)$ raised to the power 00. Since each component of $f(x)$ is a Boolean indicator for the presence or absence of a vocabulary item at a given position, we can rewrite it with Iverson brackets:
@@
In general, since the coordinates are all Booleans in $\{0,1\}$, multiplication corresponds to logical conjunction and expectation corresponds to probability:
@@ Appendix E Sampling from maximum entropy distributions with complex constraints
%**** arxiv.tex Line 800 ****L = L[…, None, :].clamp_min(0.0)
@@
%**** arxiv.tex Line 825 ****opt = optim.LBFGS([z], line_search_fn=”strong_wolfe”, max_iter=max_iter)
@@ Appendix F Computational requirements
At the scales of the datasets we use in this study, both maximum entropy hypercube-constrained sampling and Gaussian optimal transport are extremely cheap to run. In the most expensive configuration (generating around 200K $64\times 64$ CIFARNet images), the optimization loop takes roughly 65 seconds on a single Nvidia L40 GPU, while requiring approximately 29 gigabytes of GPU memory. Based on hourly pricing of \$1.10 per hour from vast.ai, this will cost around \$0.02 to generate a full set of synthetic CIFARNet images, with all other novel methods described in this paper requiring fewer computational resources than that.
@@
This means our methods can quickly become computationally infeasible with larger image sizes, which is why we limit ourselves to at most $64\times 64$ images, with memory acting as the key bottleneck for larger images.

### v2

@@ 2 Theory and Methods
where $\alpha$ is a multi-index, or a $d$-tuple assigning an integer to each coordinate of $\mathbf{x}$. Recall that taking a vector to the power of a multi-index denotes a product of the components of the vector, where each component of the index indicates the multiplicity: e.g. if $\alpha=(1,4,6)$, the expression $(\mathbf{x}-\boldsymbol{\mu})^{\alpha}$ denotes the product $(x_{1}-\mu_{1})(x_{2}-\mu_{2})^{4}(x_{3}-\mu_{3})^{6}$. Similarly, ($\partial^{\alpha}\mathcal{L}$) is shorthand for the mixed partial derivative $\frac{\partial^{11}}{\partial x_{1}\partial x_{2}^{4}\partial x_{3}^{6}}\mathcal{L}$. The factorial $\alpha!$ denotes the product of the factorials of the components: 1!×4!×6!=172801$\times$ 4$\times$ 6!=172801 ! × 4 ! × 6 ! = 17280.
@@ 2.3 Optimal Transport
Figure 4: Rows 1-3 show how Gaussian optimal transport affects the example CIFAR-10 airplane, bird and truck images. Each row starts with the original unedited image on the left, with each subsequent column showing the effects of editing that image’s first two moments to match the class-conditional distributions of a particular target class (top).
@@
Given $k$ image classes, each containing tensors of shape $C\times H\times W$, we unroll the tensors into vectors of size $CHW$, then compute their means and covariance matrices, [^7] Because this is a high-dimensional covariance matrix with dimension only 1-3 times smaller than the sample size, we apply the asymptotically optimal linear shrinkage method proposed by Bodnar et al. (2014) to improve our estimate of the population covariance and increase numerical stability. and plug these statistics into Eq. 3 to get the $k(k-1)$ optimal transport maps from each class to every other class.
@@ 2.4 Maximum Entropy Sampling
One problem with using Gaussians to generate synthetic images is that natural images are constrained to a hypercube: RGB pixel intensities are in the range [0,255][0,255][ 0 , 255 ], but nonsingular Gaussian distributions assign positive probability density to all of $\mathbb{R}^{d}$, so that a typical sample will often lie outside the hypercube of natural images. [^9] Strictly speaking, Gaussians also violate the assumption of compact support that we made earlier. In high dimension, though, almost all the probability mass of a Gaussian is contained in the typical set, a compact region near the boundary of an ellipsoid surrounding the mean (Carpenter, 2017). We might expect neural networks to quickly adapt to such a simple box constraint on the support, so we would ideally like to subject our synthetic images to this constraint.
@@ 2.5 Discrete Domains
[…]
@@ 3 Image Classification
Because Gaussian optimal transport (Sec. 2.3) requires $O(d^{3})$ compute and $O(d^{2})$ memory, [^12] The covariance matrix has $d^{2}$ elements, where $d$ is the number of pixels, so it is actually $O(n^{4})$ in the height or width dimension of the image. We also ran into a software limitation in early experiments where NumPy and PyTorch eigensolvers would crash when fed the very large covariance matrices produced by high-resolution image datasets; see PyTorch issue #92141 for discussion. we focus on datasets with $\times 32$ 32 or $\times 64$ 64 resolution images for our primary experiments. Specifically, we examine the popular image classification datasets CIFAR-10 (Krizhevsky et al., 2009), Fashion MNIST (Xiao et al., 2017), MNIST (LeCun et al., 1998), and SVHN (Netzer et al., 2011).
We also build a new image classification dataset, CIFARNet, consisting of 200K images at $\times 64$ 64 resolution sampled from ImageNet-21K, using ten coarse-grained classes that roughly match those of CIFAR-10. The larger number of images per class (20K) allows us to get a good estimate of the class-conditional covariance matrices needed for Gaussian optimal transport, which at this resolution contain $(3\times 64\times 64)^{2}\approx 1.5\times 10^{8}$ entries each. See Appendix B for more details on CIFARNet.
@@ 5 Conclusion
5.1 Acknowledgements
We are thankful to Open Philanthropy for funding this work. We also thank New Science and Stability AI for providing computing resources.
@@ Appendix B CIFARNet dataset
We ensured class balance by randomly sampling 20K images from each synset. Images were directly resized to $\times 64$ 64 resolution without center cropping.
@@ Appendix D Discrete domain proofs
where $f(x)_{1}^{0}$ denotes the first component of $f(x)$ raised to the power 0. Since each component of $f(x)$ is a Boolean indicator for the presence or absence of a vocabulary item at a given position, we can rewrite it with Iverson brackets:
@@
[…]
@@ Appendix E Sampling from maximum entropy distributions with complex constraints
L = L[…, None, :].clamp_min(0.0)
@@
opt = optim.LBFGS([z], line_search_fn=”strong_wolfe”, max_iter=max_iter)
@@ Appendix F Computational requirements
At the scales of the datasets we use in this study, both maximum entropy hypercube-constrained sampling and Gaussian optimal transport are extremely cheap to run. In the most expensive configuration (generating around 200K $\times 64$ 64 CIFARNet images), the optimization loop takes roughly 65 seconds on a single Nvidia L40 GPU, while requiring approximately 29 gigabytes of GPU memory. Based on hourly pricing of \$1.10 per hour from vast.ai, this will cost around \$0.02 to generate a full set of synthetic CIFARNet images, with all other novel methods described in this paper requiring fewer computational resources than that.
@@
This means our methods can quickly become computationally infeasible with larger image sizes, which is why we limit ourselves to at most $\times 64$ 64 images, with memory acting as the key bottleneck for larger images.